OpenStudy (andijo76):

Evaluate 12C8 (Points : 2) 24 1680 11880 495

jimthompson5910 (jim_thompson5910):

n C r = (n!)/(r!(n-r)!) 12 C 8 = (12!)/(8!*(12-8)!) 12 C 8 = (12!)/(8!*4!) 12 C 8 = (12*11*10*9*8!)/(8!*4!) 12 C 8 = (12*11*10*9)/(4!) 12 C 8 = (12*11*10*9)/(4*3*2*1) 12 C 8 = (11880)/(24) 12 C 8 = 495 So the answer is choice D

OpenStudy (andijo76):

cool thanks

jimthompson5910 (jim_thompson5910):

yw